# UniMove: A Unified Model for Multi-city Human Mobility Prediction

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2025/11
@@ 1. Introduction
Human mobility (Luca et al., 2021), which record the movement patterns of individuals across time and space, have gained substantial importance in numerous fields, including urban planning (Zheng et al., 2023a; Su et al., 2025; Zheng et al., 2023b; Su et al., 2024), transportation management (Zheng et al., 2024; Yuan et al., 2024a, b), mobile network optimization (Chai et al., 2025; Sheng et al., 2025) and personalized services (Lian et al., 2015; Zheng et al., 2010; Yuan et al., 2023b; Zheng and Xie, 2011). Human mobility presents unique challenges in modeling due to their inherent randomness, non-uniform time intervals, and complex patterns (Yuan et al., 2025a). Unlike vehicle trajectories (Zhu et al., 2023), which typically follow more predictable and regular paths with time intervals on the order of seconds, human mobility data is more irregular and context-dependent.
@@
To effectively tackle such data and harness the wealth of mobility information, researchers often treat a mobility trajectory as a sequence of location tokens, with each city represented by a set of unique tokens, similar to a vocabulary in language processing. These tokens are typically discretized using grid-based representations (Long et al., 2024; Yuan et al., 2025b), administrative divisions (Yuan et al., 2025a), or points of interest (POIs) (Feng et al., 2018; Shao et al., 2024), providing a structured approach to analyzing movement patterns. Then, advanced AI models, such as RNNs (Feng et al., 2018; Yuan et al., 2022), transformers (Si et al., 2023; Long et al., 2024), and other deep learning architectures (Yuan et al., 2023a; Shao et al., 2024; Zhu et al., 2023), can be utilized to learn the underlying patterns and dynamics of mobility trajectories. However, these tokens do not overlap across cities due to the unique geographic coverage of each urban area. As a result, practitioners often need to train separate models for different geographic regions, which limits generalizability and practical applicability. While some exceptions exist, such as transfer learning approaches (Zhu et al., 2024a; He et al., 2024a; Xu and Guan, 2024; Wang et al., 2024), they either rely on vehicle trajectories (where location representations, such as numerical latitude and longitude, are transferable) (Zhu et al., 2024a) or just use more tokens (He et al., 2024a) or separate encoders (Wang et al., 2024) to address geographic differences. Some methods, such as ST-MoE-BERT  (He et al., 2024b), force the grid cells of different cities into the same number of classes to process mobility data from different cities with a shared encoder, which severely eroding the distinct spatial identities of individual cities.
@@ 2.1. Human Mobility Prediction
Human mobility prediction has been extensively studied using various modeling techniques, each aiming to capture the intricate dynamics of human movement. Real-world applications of human mobility prediction span a variety of domains, including smart cities (Wang et al., 2020; Yuan et al., 2024d), traffic management (Yuan et al., 2024a, b), personalized recommendations (Lian et al., 2015; Zheng et al., 2010; Zheng and Xie, 2011), and public health (e.g., predicting the spread of diseases) (Nouvellet et al., 2021; Han et al., 2021; Chang et al., 2021). Early methods relied heavily on probabilistic models, such as the Markov model (Chen et al., 2014; Gambs et al., 2012) and EPR-based models (Song et al., 2010; Jiang et al., 2016), where mobility patterns are derived from the transition probabilities between discrete locations. These approaches, while effective in some contexts, often fail to capture the complex, dynamic nature of human behavior (Yuan et al., 2025a; Feng et al., 2018), especially when dealing with non-uniform, irregular time intervals between events. Recently, deep learning models have significantly advanced mobility prediction. Models such as recurrent neural networks (RNNs) (Chen et al., 2020), attention networks (Yuan et al., 2023a; Feng et al., 2018), Transformers (Corrias et al., 2023; Si et al., 2023), Graph Neural Networks (GNNs) (Sun et al., 2021; Terroso-Sáenz and Muñoz, 2022), diffusion models (Long et al., 2025; Zhu et al., 2023; Long et al., 2024) have been employed to capture the temporal dependencies inherent in mobility data. However, these models typically require training with large-scale data. Since mobility patterns are highly localized and influenced by unique geographic, cultural, and infrastructural factors, they necessitate training separate models for each city’s data.
@@ 2.2. Location Encoding
Location encoding (Mai et al., 2022) plays a crucial role in human mobility prediction by transforming spatial data into a form that can be processed by deep neural networks. Location encoders can be divided into two main categories: continuous location encoding and discrete location encoding. For continuous encoding, locations are treated as continuous values, often represented by continuous coordinates (e.g., latitude, longitude) (Zhu et al., 2024b; Zhang et al., 2024; Yuan et al., 2022, 2024c). These continuous values can be directly embedded into the model through continuous embeddings (Cao and Li, 2021; Zhu et al., 2024b; Zhang et al., 2024) or positional encodings (Yuan et al., 2022, 2023a), which are typically processed via neural networks to capture spatial relationships in a continuous space. In discrete location encoding, the continuous spatial domain is often divided into predefined locations, such as Points of Interest (POI) (Feng et al., 2018; Shao et al., 2024), Areas of Interest (AOI) (Zhang et al., 2023), administrative districtc (Yuan et al., 2025a), or grids (Long et al., 2024; Yuan et al., 2025a). Each of these locations is assigned a discrete identifier, and the locations are then represented using techniques like one-hot encoding (Brownlee, 2020). In this way, each location is transformed into a unique token, much like words in language models. Discrete encoding simplifies the representation, making it more computationally efficient while still allowing the model to process and distinguish between different locations effectively.
@@ 2.3. Cross-City Modeling of Human Mobility
Cross-city modeling of human mobility aims to leverage knowledge and patterns learned from one city to improve mobility modeling in another, addressing the challenges posed by data scarcity and the unique characteristics of individual cities (He et al., 2020; Xu and Guan, 2024; Wang et al., 2024; Jiang et al., 2021; Gupta and Bedathur, 2022). For instance, He et al. (He et al., 2020) propose a top-down paradigm that utilizes abundant annotated data to plan mobility for a new city. Additionally, meta-learning frameworks (Vilalta and Drissi, 2002) have been developed to enable models to adapt quickly to new cities with limited data (Gupta and Bedathur, 2022; Wang et al., 2024). However, existing approaches still face significant limitations. For example, they often fail to address the challenge of using a single shared model (with identical parameters) across different cities. Current methods either rely on separate location embedding layers for each city (Wang et al., 2024) or are restricted to handling vehicle trajectories with numerical coordinate values (Zhu et al., 2024a, b). Unlike vehicle trajectories, which primarily rely on numerical GPS coordinates (e.g., latitude and longitude), human mobility trajectories are inherently richer in semantics, encompassing contextual information such as PoIs, activities, and urban infrastructure. This requires to address the tokenization of locations across cities, ensuring that models can effectively generalize and adapt to diverse urban environments.
@@ 4.3. Deep & Cross Net
Location Tower introduces the Deep & Cross Net(DCN) to further capture features from different locations. It consists of two main parts: the Cross layer and the Deep layer. The design of the Cross layer is inspired by  (Wang et al., 2017), aiming to capture high-order feature interactions through feature crossing. Suppose the input feature vector is location embedding $\mathbf{E_{l}}$, with a dimension of $d$. The Cross layer calculates as follows:
@@ 5.1.1. Datasets
5.1.2. Metrics
5.1.3. Baselines
  • Markov (Gambs et al., 2012):The Markov model serves as a statistical framework designed to depict how states evolve over time. It predicts future locations by computing the probabilities of transitions.
  • LSTM (Kong and Wu, 2018):LSTM networks excel at processing sequential data and are adept at capturing long-term dependencies, making them naturally suitable for location prediction tasks.
  • Transformer (Vaswani, 2017):Transformer is a deep learning architecture based on the self-attention mechanism, capable of efficiently processing sequential data and capturing long-range dependencies.
  • DeepMove (Feng et al., 2018):This approach is particularly effective in addressing the challenges of predicting human movement patterns, which are influenced by various factors such as time, location, and personal preferences
  • TrajBert (Si et al., 2023):TrajBERT is a trajectory recovery method based on BERT, which recovers implicit sparse trajectories through spatial-temporal refinement.
  • STAN (Luo et al., 2021): STAN employs a spatio-temporal bi-attention architecture to capture non-adjacent location correlations and non-consecutive visit dependencies.
  • GETNext (Yang et al., 2022): GETNext integrates Graph Convolutional Networks (GCN) and Transformer to model global POI transition patterns via a trajectory flow map. It fuses spatio-temporal contexts with time-aware category embeddings for next POI recommendation.
  • CTLE (Lin et al., 2021): The Context and Time aware Location Embedding (CTLE) model dynamically generates context-specific location embeddings using bidirectional Transformers and temporal encoding to capture multi-functional properties of locations under varying contexts and temporal dynamics.
5.1.4. Implementation Details
5.2. Overall Performance

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6,2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762780isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Machine learningccs: Information systems Mobile information processing systems11footnotetext: Equal contribution22footnotetext: Corresponding authors
@@ 1. Introduction
Human mobility (luca2021survey), which record the movement patterns of individuals across time and space, have gained substantial importance in numerous fields, including urban planning (zheng2023spatial; sureinforcement; zheng2023road; su2024metrognn), transportation management (zheng2024survey; yuan2024unist; yuan2024foundation), mobile network optimization (chai2025mobiworld; sheng2025unveiling) and personalized services (lian2015content; zheng2010collaborative; yuan2023learning; zheng2011learning). Human mobility presents unique challenges in modeling due to their inherent randomness, non-uniform time intervals, and complex patterns (yuan2025learning). Unlike vehicle trajectories (zhu2023difftraj), which typically follow more predictable and regular paths with time intervals on the order of seconds, human mobility data is more irregular and context-dependent.
To effectively tackle such data and harness the wealth of mobility information, researchers often treat a mobility trajectory as a sequence of location tokens, with each city represented by a set of unique tokens, similar to a vocabulary in language processing. These tokens are typically discretized using grid-based representations (long2024universal; yuan2025worldmove), administrative divisions (yuan2025learning), or points of interest (POIs) (feng2018deepmove; shao2024beyond), providing a structured approach to analyzing movement patterns. Then, advanced AI models, such as RNNs (feng2018deepmove; yuan2022activity), transformers (si2023trajbert; long2024universal), and other deep learning architectures (yuan2023spatio; shao2024beyond; zhu2023difftraj), can be utilized to learn the underlying patterns and dynamics of mobility trajectories. However, these tokens do not overlap across cities due to the unique geographic coverage of each urban area. As a result, practitioners often need to train separate models for different geographic regions, which limits generalizability and practical applicability. While some exceptions exist, such as transfer learning approaches (zhu2024controltraj; he2024st; xu2024crosspred; wang2024cola), they either rely on vehicle trajectories (where location representations, such as numerical latitude and longitude, are transferable) (zhu2024controltraj) or just use more tokens (he2024st) or separate encoders (wang2024cola) to address geographic differences. Some methods, such as ST-MoE-BERT (stbertmoe), force the grid cells of different cities into the same number of classes to process mobility data from different cities with a shared encoder, which severely erodes the distinct spatial identities of individual cities.
@@ 2.1. Human Mobility Prediction
Human mobility prediction has been extensively studied using various modeling techniques, each aiming to capture the intricate dynamics of human movement. Real-world applications of human mobility prediction span a variety of domains, including smart cities (wang2020review; yuan2024urbandit), traffic management (yuan2024unist; yuan2024foundation), personalized recommendations (lian2015content; zheng2010collaborative; zheng2011learning), and public health (e.g., predicting the spread of diseases) (nouvellet2021reduction; han2021will; chang2021mobility). Early methods relied heavily on probabilistic models, such as the Markov model (chen2014nlpmm; gambs2012next) and EPR-based models (song2010modelling; jiang2016timegeo), where mobility patterns are derived from the transition probabilities between discrete locations. These approaches, while effective in some contexts, often fail to capture the complex, dynamic nature of human behavior (yuan2025learning; feng2018deepmove), especially when dealing with non-uniform, irregular time intervals between events. Recently, deep learning models have significantly advanced mobility prediction. Models such as recurrent neural networks (RNNs) (chen2020predicting), attention networks (yuan2023spatio; feng2018deepmove), Transformers (corrias2023exploring; si2023trajbert), Graph Neural Networks (GNNs) (sun2021periodicmove; terroso2022nation), diffusion models (long2025one; zhu2023difftraj; long2024universal) have been employed to capture the temporal dependencies inherent in mobility data. However, these models typically require training with large-scale data. Since mobility patterns are highly localized and influenced by unique geographic, cultural, and infrastructural factors, they necessitate training separate models for each city’s data.
@@ 2.2. Location Encoding
Location encoding (mai2022review) plays a crucial role in human mobility prediction by transforming spatial data into a form that can be processed by deep neural networks. Location encoders can be divided into two main categories: continuous location encoding and discrete location encoding. For continuous encoding, locations are treated as continuous values, often represented by continuous coordinates (e.g., latitude, longitude) (zhu2024unitraj; zhang2024noise; yuan2022activity; yuan2024generating). These continuous values can be directly embedded into the model through continuous embeddings (cao2021generating; zhu2024unitraj; zhang2024noise) or positional encodings (yuan2022activity; yuan2023spatio), which are typically processed via neural networks to capture spatial relationships in a continuous space. In discrete location encoding, the continuous spatial domain is often divided into predefined locations, such as Points of Interest (POI) (feng2018deepmove; shao2024beyond), Areas of Interest (AOI) (zhang2023city), administrative districtc (yuan2025learning), or grids (long2024universal; yuan2025learning). Each of these locations is assigned a discrete identifier, and the locations are then represented using techniques like one-hot encoding (brownlee2020ordinal). In this way, each location is transformed into a unique token, much like words in language models. Discrete encoding simplifies the representation, making it more computationally efficient while still allowing the model to process and distinguish between different locations effectively.
@@ 2.3. Cross-City Modeling of Human Mobility
Cross-city modeling of human mobility aims to leverage knowledge and patterns learned from one city to improve mobility modeling in another, addressing the challenges posed by data scarcity and the unique characteristics of individual cities (he2020human; xu2024crosspred; wang2024cola; jiang2021transfer; gupta2022doing). For instance, He et al. (he2020human) propose a top-down paradigm that utilizes abundant annotated data to plan mobility for a new city. Additionally, meta-learning frameworks (vilalta2002perspective) have been developed to enable models to adapt quickly to new cities with limited data (gupta2022doing; wang2024cola). However, existing approaches still face significant limitations. For example, they often fail to address the challenge of using a single shared model (with identical parameters) across different cities. Current methods either rely on separate location embedding layers for each city (wang2024cola) or are restricted to handling vehicle trajectories with numerical coordinate values (zhu2024controltraj; zhu2024unitraj). Unlike vehicle trajectories, which primarily rely on numerical GPS coordinates (e.g., latitude and longitude), human mobility trajectories are inherently richer in semantics, encompassing contextual information such as PoIs, activities, and urban infrastructure. This requires to address the tokenization of locations across cities, ensuring that models can effectively generalize and adapt to diverse urban environments.
@@ 4.3. Deep & Cross Net
Location Tower introduces the Deep & Cross Net(DCN) to further capture features from different locations. It consists of two main parts: the Cross layer and the Deep layer. The design of the Cross layer is inspired by  (wang2017deep), aiming to capture high-order feature interactions through feature crossing. Suppose the input feature vector is location embedding $\mathbf{E_{l}}$, with a dimension of $d$. The Cross layer calculates as follows:
@@ 5.1.1. Datasets
5.1.2. Metrics
5.1.3. Baselines
  • Markov (gambs2012next):The Markov model serves as a statistical framework designed to depict how states evolve over time. It predicts future locations by computing the probabilities of transitions.
  • LSTM (kong2018hst):LSTM networks excel at processing sequential data and are adept at capturing long-term dependencies, making them naturally suitable for location prediction tasks.
  • Transformer (vaswani2017attention):Transformer is a deep learning architecture based on the self-attention mechanism, capable of efficiently processing sequential data and capturing long-range dependencies.
  • DeepMove (feng2018deepmove):This approach is particularly effective in addressing the challenges of predicting human movement patterns, which are influenced by various factors such as time, location, and personal preferences
  • TrajBert (si2023trajbert):TrajBERT is a trajectory recovery method based on BERT, which recovers implicit sparse trajectories through spatial-temporal refinement.
  • STAN (luo2021stan): STAN employs a spatio-temporal bi-attention architecture to capture non-adjacent location correlations and non-consecutive visit dependencies.
  • GETNext (yang2022getnext): GETNext integrates Graph Convolutional Networks (GCN) and Transformer to model global POI transition patterns via a trajectory flow map. It fuses spatio-temporal contexts with time-aware category embeddings for next POI recommendation.
  • CTLE (lin2021pre): The Context and Time aware Location Embedding (CTLE) model dynamically generates context-specific location embeddings using bidirectional Transformers and temporal encoding to capture multi-functional properties of locations under varying contexts and temporal dynamics.
5.1.4. Implementation Details
5.2. Overall Performance
@@ 6. Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under 62476152 and U24B20180.